How to find the coordinate of p and q if the circle cuts the x-axis at the points p and q in 2x^2 +2y^2 -8x +5y -10=0

1
Expert's answer
2022-02-14T13:04:58-0500

Since the circle cuts the x-axis then "y=0".

Substituting "y=0" into the given equation of the circle, "2x^2 +2y^2 -8x +5y -10=0", we have

"2x^2-8x-10=0\\
\Rightarrow x^2-4x-5=0\\
\Rightarrow (x-5)(x+1)=0\Rightarrow x=-1,5"

Thus, without loss of generality, the circle cuts the x-axis at P"(-1,0)" and Q"(5,0)"